# BASE-Q: Bias and Asymmetric Scaling Enhanced Rotational Quantization for Large Language Models

### v1

@@ 5.4 Resource and accuracy trade-offs
Recent studies such as FlatQuantsun2024flatquant employ layer-wise learned online matrix transforms to improve the performance of quantized linear layers, at the cost of increased inference overhead. In contrast, BASE-Q applies rotations $\boldsymbol{\mathit{R}}_{res}$ that can be fused into the weights, eliminating the need for the extra computation. As shown in Figure 7, BASE-Q consistently achieves Pareto-optimal trade-offs between accuracy and inference overhead, outperforming Quarot at comparable cost, while FlatQuant suffers from significant computation overhead. Furthermore, FlatQuant suffers from significant instability on Qwen2.5-3B, where the quantized model collapses. Figure 8 further demonstrates the lower GPU memory usage of BASE-Q under quantization process. Importantly, our method is fully orthogonal to other advanced methods and can flexibly combine with them, allowing practitioners to efficiently adapt to different deployment and training requirements.
@@ Appendix A Full quantization results
We present comprehensive quantization results in this section, including perplexity on WikiText2 and zero-shot accuracy on nine evaluation datasets. All baseline results, including QuaRot, SpinQuant, and OSTQuant, are reproduced using their official open-source implementations, with necessary modifications to support the Qwen model, which differs from LLaMA mainly by including attention bias. As the official OSTQuant repository does not support FSDP (Fully Sharded Data Parallel) training, its results are limited to models with up to 14B parameters, while larger models, such as the 70B model, encounter out-of-memory (OOM) issues. Results for the Llama2 series are summarized in Appendix A, those for the Llama3 series are in Appendix A, and for the Qwen2.5 series in Appendix A.
@@ Appendix B Visualization of activation distribution
We provide visualization results for the input activations of the multi-head attention blocks and MLP blocks at the 1st, 11th, and 31st layers of Qwen2.5-3B, Llama2-7B, and Llama3-8B. For each selected layer, we illustrate the distributions of activations under various rotation strategies. We observe that inter-channel mean misalignment consistently occurs in different layers and across all evaluated models, and that learned rotations are insufficient to eliminate this issue. As discussed in Section 3.2, this misalignment is a significant source of quantization error, which can be effectively mitigated by our proposed bias correction technique.
@@ Appendix C Composition of rounding errors
This section presents visualizations of the composition of rounding errors, calculated according to Equations 5 and 6, across various models.

### v2

@@ 5.4 Resource and accuracy trade-offs
Recent studies such as FlatQuantsun2024flatquant employ layer-wise learned online matrix transforms to improve the performance of quantized linear layers, at the cost of increased inference overhead. In contrast, BASE-Q applies rotations $\boldsymbol{\mathit{R}}_{res}$ that can be fused into the weights, eliminating the need for the extra computation. As shown in Figure˜7, BASE-Q consistently achieves Pareto-optimal trade-offs between accuracy and inference overhead, outperforming Quarot at comparable cost, while FlatQuant suffers from significant computation overhead. Furthermore, FlatQuant suffers from significant instability on Qwen2.5-3B, where the quantized model collapses. Figure˜8 further demonstrates the lower GPU memory usage of BASE-Q under quantization process. Importantly, our method is fully orthogonal to other advanced methods and can flexibly combine with them, allowing practitioners to efficiently adapt to different deployment and training requirements.
@@ Appendix A Full quantization results
We present comprehensive quantization results in this section, including perplexity on WikiText2 and zero-shot accuracy on nine evaluation datasets. All baseline results, including QuaRot, SpinQuant, and OSTQuant, are reproduced using their official open-source implementations, with necessary modifications to support the Qwen model, which differs from LLaMA mainly by including attention bias. As the official OSTQuant repository does not support FSDP (Fully Sharded Data Parallel) training, its results are limited to models with up to 14B parameters, while larger models, such as the 70B model, encounter out-of-memory (OOM) issues. Results for the Llama2 series are summarized in Table˜4, those for the Llama3 series are in Table˜5, and for the Qwen2.5 series in Table˜6.
@@ Appendix B Visualization of activation distribution
We provide visualization results for the input activations of the multi-head attention blocks and MLP blocks at the 1st, 11th, and 31st layers of Qwen2.5-3B, Llama2-7B, and Llama3-8B. For each selected layer, we illustrate the distributions of activations under various rotation strategies. We observe that inter-channel mean misalignment consistently occurs in different layers and across all evaluated models, and that learned rotations are insufficient to eliminate this issue. As discussed in Section˜3.2, this misalignment is a significant source of quantization error, which can be effectively mitigated by our proposed bias correction technique.
@@ Appendix C Composition of rounding errors
This section presents visualizations of the composition of rounding errors, calculated according to Equations˜5 and 6, across various models.